# Making Teams and Influencing Agents: Efficiently Coordinating Decision Trees for Interpretable Multi-Agent Reinforcement Learning

### v1

@@ 5.4 Hyperparameter Sensitivity
Figure 4: Sensitivity of HYDRAVIPER to two hyperparameters, DT depth and the UCB scaling constant $c$, on the cooperative navigation (CN) environment. Lower rewards are better. The default value is 4 for both hyperparameters.
To understand the effects of HYDRAVIPER’s hyperparameters on its performance, we conduct experiments to vary the depth of the DT students, and the scaling constant $c$ for UCB policy selection (Section 4.3), on the cooperative navigation environment. We choose this environment due to its relatively low level of randomness. For these experiments, we use HYDRAVIPER at the low budget level (500 training/1 500 validation rollouts) as the baseline algorithm, and fix all hyperparameters other than those of interest.
By default, we use a DT depth of 4; our results show that DTs of this depth provide a good tradeoff between expressiveness and computational efficiency. As shown on the left of Figure 4, depth-4 DTs outperform depth-2 and depth-3 DTs on cooperative navigation. This is an intuitive result; the optimal agent policy for this environment cannot be represented with such shallow DTs, as they must condition on the positions of the other agents and the landmarks. However, we find that depth-4 DTs also marginally outperform depth-5 DTs. This same pattern exists in all of the environments that we use for evaluation. We hypothesise that the amount of data collected by HYDRAVIPER at the low budget level is insufficient to coordinate between depth-5 DTs.
Our default value for $c$ is also 4. As we outlined in Corollary 1 in Section 4.3, the value that theoretically allows logarithmic regret to be achieved is based on the range of mean returns in the environment. Since the agents navigate in a $2\times 2$ square environment, the maximum distance of an agent to a target is $2\sqrt{2}$. The reward in this environment is the negation of the minimum agent distance to each landmark, plus a penalty of -1 for each agent that the ego agent collides with. Therefore, the maximum possible penalty is $\Delta=3\cdot 2\sqrt{2}+2$, which requires $c\approx 219.88$ to achieve the guarantee of Corollary 1. However, as shown on the right of Figure 4, $c=4$ empirically performs best; $c=8$ and $c=16$ are already excessively conservative given the low randomness in the environment. Since the agents are already trained, we hypothesise that the potential range of returns is less useful in practice for finding a good policy profile than the typical range of returns. This can be approximated by performing expert rollouts, as we described in Section 4.3.
Sensitivity results for the three other environments are shown in Appendix D. The best DT depths for these environments are all 4, as with cooperative navigation, while the best tested values of $c$ for physical deception and the Cologne corridor are respectively 2 and 16.
Figure 5: Ablation of HYDRAVIPER on the cooperative navigation (CN, left) and Cologne corridor (CC, right) environments. Lower rewards are better.
@@ 5.5 Ablation
Lastly, to understand which components of HYDRAVIPER are responsible for its success, we conduct an ablation study for HYDRAVIPER at two budget levels in the cooperative navigation (CN) and Cologne corridor (CC) environments. For the high budget level, we use 5 000 training/5 000 validation rollouts for CN, and 1 000 training/1 000 validation rollouts for CC; for the low budget level, we use 500 training/1 500 validation rollouts for CN, and 100 training/100 validation rollouts for CC. We compare HYDRAVIPER’s centralised-$Q$ resampling with IVIPER’s independent resampling (HYDRAVIPER - CQ). Also, we study the impact of removing adaptive training budget allocation (HYDRAVIPER - TR-A) and UCB-based validation budget allocation (HYDRAVIPER - VR-UCB). Results for physical deception and the Ingolstadt corridor are shown in Appendix E.
Figure 5 shows the results of our ablation study. On both environments and at both budget levels, centralised-$Q$ resampling outperforms the IVIPER resampling scheme, although the gap in performance is less pronounced for the Cologne corridor due to environmental randomness. This result suggests that sampling the training dataset independently for each agent, instead of according to team performance, is insufficient for achieving coordinated behaviour in the resulting students. Meanwhile, removing the budget allocation methods degrades the performance of HYDRAVIPER. Having either one of the budget allocation methods is generally sufficient to improve HYDRAVIPER’s reward, except in one case: for cooperative navigation at the low budget level, HYDRAVIPER performs worse when only one budget allocation mechanism is present. Meanwhile, for the Cologne corridor at the low budget level, the variance in HYDRAVIPER’s reward is large both when only centralised-$Q$ resampling is present, and also when it is removed. These results suggest that the primary benefit of the two rollout budget allocation mechanisms is to stabilise HYDRAVIPER’s learning process, especially in the low budget setting when extracting the most information from each rollout is critical.
@@ 6 Conclusion and Future Work
In this work, we introduced a new DT-based interpretable MARL method, HYDRAVIPER. HYDRAVIPER addresses several limitations of prior multi-agent methods that follow the VIPER framework: (1) it improves performance by using a joint dataset resampling scheme based on team $Q$-values, and (2) it improves computational efficiency by adaptively allocating fixed budgets of environment interactions for training and validation, as well as by dividing agents into jointly-trained teams. Based on experiments in benchmark environments for multi-agent coordination and traffic signal control, we showed that HYDRAVIPER achieves performance comparable with MAVIPER (a centralised method) and even neural network experts, all within a runtime less than IVIPER (a decentralised method). We also demonstrated HYDRAVIPER’s sample efficiency in its ability to retain a similar level of performance using a fraction of the environment interactions.
Through our experiments in the Ingolstadt corridor environment, we scaled up the VIPER framework to seven agents. To our knowledge, this is the largest team of coordinated agents to which interpretable MARL has been applied thus far. However, environments based on real-world domains can have many more agents than the environments that we studied. For instance, the review of Noaeen et al. (2022) showed that TSC environments of dozens or even hundreds of agents are used in the RL literature. At the extreme, Chen et al. (2020) used parameter-shared MPLight policies as controller agents for an extremely large simulation of 2 510 traffic lights. Our agent clustering approach shows promise in scaling up to larger environments while retaining performance comparable to that of expert policies. In scaling up, other methods of assigning edge weights $e_{ij}$ to agent pairs could be developed to leverage structure present in MARL environments, including weights based on $Q$-values or on observation-action trajectories. We envision that the flexibility of the HYDRAVIPER framework will allow it to adapt to characteristics of different MARL environments while maintaining Pareto optimality in the performance-computational efficiency tradeoff.
@@ Appendix A Proof of Resampling Theorem
None of $\mathbf{o}_{k}$, $\pi^{*}_{-\ell}$, $\bar{V}^{\pi_{\ell}^{*}}$, or $\bar{Q}^{\pi_{\ell}^{*}}$ changed between iteration $m$ and $m+1$, since HYDRAVIPER uses experts and expert value functions to compute the weights. This contradicts Equation 2. Thus, if $p_{\ell k}^{(m)}=0$, then $p_{\ell k}^{(m+1)}=0$. ∎
@@
Appendix B Full Performance Results
@@ Appendix C Full Runtime Results
Figure 6: Runtime decomposition for IVIPER, MAVIPER, and HYDRAVIPER (with full environment interaction budget) on the two multi-agent particle world environments, physical deception and cooperative navigation. Error bars show 95% confidence intervals based on 10 random seeds. Runtime measurement was performed following the methodology in Section 5.2.
Figure 7: Runtime decomposition for IVIPER, MAVIPER, and HYDRAVIPER (with full environment interaction budget) on the two traffic signal control environments, Cologne corridor and Ingolstadt corridor. Error bars show 95% confidence intervals based on 10 random seeds. Runtime measurement was performed following the methodology in Section 5.2.
Appendix D Full Hyperparameter Sensitivity Results
In Figure 8, Figure 9, and Figure 10, we show the performance of HYDRAVIPER in the low-budget setting as two hyperparameters — DT depth and the UCB scaling constant $c$ — are varied for the physical deception, cooperative navigation, and Ingolstadt corridor environments.
@@ Appendix D Full Hyperparameter Sensitivity Results
Figure 9: Sensitivity of HYDRAVIPER to two hyperparameters, DT depth and the UCB scaling constant $c$, on the cooperative navigation (CN) environment. Lower rewards are better. The default value is 4 for both hyperparameters.
@@
Appendix E Full Ablation Results
In Figure 11, we show the performance of HYDRAVIPER on the physical deception (left) and Ingolstadt corridor (right) environments, in an ablation study where each of the algorithm’s components — centralised-$Q$ resampling (C-Q), adaptive training budget allocation (TR-A), and UCB-based validation budget allocation (VR-UCB) — is removed in turn.
Figure 11: Ablation of HYDRAVIPER on the physical deception (PD, left, higher rewards are better) and Ingolstadt corridor (IC, right, lower rewards are better) environments.

### v2

@@ 5.4 Hyperparameter Sensitivity
To understand the effects of HYDRAVIPER’s hyperparameters on its performance, we conduct experiments to vary the depth of the DT students, and the scaling constant $c$ for UCB policy selection (Section 4.3), on the cooperative navigation environment. We choose this environment due to its relatively low level of randomness. For these experiments, we use HYDRAVIPER at the low budget level (500 training/1 500 validation rollouts) as the baseline algorithm, and fix all hyperparameters other than those of interest. Figures for all results are shown in Appendix E.
By default, we use a DT depth of 4; our results show that DTs of this depth provide a good tradeoff between expressiveness and computational efficiency. Depth-4 DTs outperform depth-2 and depth-3 DTs on cooperative navigation. This is an intuitive result; the optimal agent policy for this environment cannot be represented with such shallow DTs, as they must condition on the positions of the other agents and the landmarks. However, we find that depth-4 DTs also marginally outperform depth-5 DTs. This same pattern exists in all of the environments that we use for evaluation. We hypothesise that the amount of data collected by HYDRAVIPER at the low budget level is insufficient to coordinate between depth-5 DTs.
Our default value for $c$ is also 4. As discussed in Appendix B, this value is smaller than would be necessary according to a theoretical analysis. Since the agents are already trained, we hypothesise that the potential range of returns is less useful in practice for finding a good policy profile than the typical range of returns. The best DT depths for the other environments are all 4, as with cooperative navigation, while the best tested values of $c$ for physical deception and the Cologne corridor are, respectively, 2 and 16.
Figure 5: Ablation on cooperative navigation (CN, L) and Cologne corridor ($C$, R). Lower rewards are better.
@@ 5.5 Ablation
Lastly, to understand which components of HYDRAVIPER are responsible for its success, we conduct an ablation study for HYDRAVIPER at two budget levels in the cooperative navigation (CN) and Cologne corridor (CC) environments. For the high budget level, we use 5 000 training/5 000 validation rollouts for CN, and 1 000 training/1 000 validation rollouts for CC; for the low budget level, we use 500 training/1 500 validation rollouts for CN, and 100 training/100 validation rollouts for CC. We compare HYDRAVIPER’s centralised-$Q$ resampling with IVIPER’s independent resampling (HYDRAVIPER - CQ). In addition, we study the impact of removing adaptive training budget allocation (HYDRAVIPER - TR-A) and UCB-based validation budget allocation (HYDRAVIPER - VR-UCB). Results for physical deception and the Ingolstadt corridor are shown in Appendix F.
Figure 5 shows our ablation results. In both environments and at both budget levels, centralised-$Q$ resampling outperforms the IVIPER resampling scheme, although the performance gap is less pronounced for the Cologne corridor due to environmental randomness. This result suggests that sampling the training dataset independently for each agent, instead of according to team performance, is insufficient to achieve coordinated behaviour in the resulting students. Meanwhile, removing the budget allocation methods degrades the performance of HYDRAVIPER. Having either one of the budget allocation methods is generally sufficient to improve HYDRAVIPER’s reward, except in one case: for cooperative navigation at the low budget level, HYDRAVIPER performs worse when only one budget allocation mechanism is present. Meanwhile, for the Cologne corridor at the low budget level, the variance in HYDRAVIPER’s reward is large both when only centralised-$Q$ resampling is present, and also when it is removed. These results suggest that the primary benefit of the two rollout budget allocation mechanisms is to stabilise HYDRAVIPER’s learning process, especially in the low budget setting when extracting the most information from each rollout is critical.
@@ 6 Conclusion and Future Work
In this work, we introduced a new DT-based interpretable MARL method, HYDRAVIPER. HYDRAVIPER addresses several limitations of prior multi-agent methods that follow the VIPER framework: (1) it improves performance by using a joint dataset resampling scheme based on team $Q$-values, and (2) it improves computational efficiency by adaptively allocating fixed budgets of environment interactions for training and validation, as well as by dividing agents into jointly-trained teams. Based on experiments in benchmark environments for multi-agent coordination and traffic signal control, we showed that HYDRAVIPER achieves performance comparable with MAVIPER (a centralised method) and even neural network experts, all with a runtime less than IVIPER (a decentralised method). We also demonstrated HYDRAVIPER’s sample efficiency in its ability to retain a similar level of performance using a fraction of the environment interactions.
Through our experiments in the Ingolstadt corridor environment, we scaled up the VIPER framework to seven agents. To our knowledge, this is the largest team of coordinated agents to which interpretable MARL has been applied so far. However, environments based on real-world domains can have many more agents than the environments that we studied. For example, the review of Noaeen et al. (2022) showed that TSC environments of dozens or even hundreds of agents are used in the RL literature. In the most extreme case, Chen et al. (2020) used parameter-shared MPLight policies as controller agents for an extremely large simulation of 2 510 traffic lights. Our agent clustering approach shows promise in scaling up to larger environments while retaining performance comparable to that of expert policies. We envision that the flexibility of the HYDRAVIPER framework will allow it to adapt to characteristics of different MARL environments while maintaining Pareto optimality in the performance-computational efficiency tradeoff.
However, our functionally grounded evaluation has not shown whether the resulting DTs are sufficient to help stakeholders better understand these policies. An application-grounded evaluation (Doshi-Velez and Kim 2017) of HYDRAVIPER that includes user studies would be necessary to assess its practical utility. Combining the general framework of HYDRAVIPER with alternative policy structures, including those based on natural language, may help make these DT-based policies more understandable and controllable.
Acknowledgments
This work was supported by the Tang Family Endowed Innovation Fund and NSF grant IIS-2046640 (CAREER). Additionally, we thank Naveen Raman, Yixuan Xu, Jingwu Tang, Matteo Pozzi, and Peter Stone for their feedback.
@@ Appendix A Proof of Resampling Theorem
None of $\mathbf{o}_{k}$, $\pi^{*}_{-\ell}$, $\bar{V}^{\pi_{\ell}^{*}}$, or $\bar{Q}^{\pi_{\ell}^{*}}$ changed between iterations $m$ and $m+1$, since HYDRAVIPER uses experts and expert value functions to compute the weights. This contradicts Equation 2. Thus, if $p_{\ell k}^{(m)}=0$, then $p_{\ell k}^{(m+1)}=0$. ∎
@@
Appendix B Application of UCB1 to HYDRAVIPER
Although UCB assumes that the arms are bounded in $[0,1]$, it can be modified in a manner equivalent to rescaling the rewards to remain in $[0,1]$. HYDRAVIPER relies on the general form of the Chernoff-Hoeffding bound:
Theorem 2.
Corollary 1.
This is the same bound as demonstrated for UCB1, and the same holds for the lower confidence bound $\mu_{\ell}^{m}-\sqrt{\frac{c\ln B}{n_{m}}}$. Overall, this choice of $c$ leads to the same $O(\log B)$ regret bound as UCB1. However, HYDRAVIPER can also be extended to use other MAB algorithms. If the mean returns of each policy profile are assumed to be normally distributed, the UCB1-Normal algorithm (Auer, Cesa-Bianchi, and Fischer 2002) could be used; it also achieves logarithmic regret. This algorithm effectively chooses $c$ to be proportional to the arms’ sample variance; the greater the variance, the wider the confidence bound. An offline (but biased) estimate of the sample variance can also be obtained by performing expert rollouts before running UCB.
How do these theoretical results apply empirically? In the cooperative navigation environment, the agents navigate in a $2\times 2$ square environment, and thus the maximum distance of an agent to a target is $2\sqrt{2}$. The reward in this environment is the negation of the minimum agent distance to each landmark, plus a penalty of -1 for each agent that the ego agent collides with. Therefore, the maximum possible penalty is $\Delta=3\cdot 2\sqrt{2}+2$, which requires $c\approx 219.88$ to achieve the guarantee of Corollary 1. However, as shown on the right of Figure 8, $c=4$ empirically performs the best; $c=8$ and $c=16$ are already excessively conservative given the low randomness in the environment. In physical deception, the reward is the $L_{2}$ distance of the adversary to the goal, minus the minimum $L_{2}$ distance of any defender to the goal. Again, the maximum distance from an agent to the goal is $\Delta=2\sqrt{2}$. Therefore, $c=16$ is required to achieve the guarantee of Corollary 1. In practice, $c=2$ still led to good performance. In the TSC environments, the reward in these environments is the total queue length (i.e. the total number of stopped vehicles in all lanes), which is effectively unbounded, but the best values empirically are $c=16$ and $c=4$.
Appendix C Full Performance Results
@@
Appendix D Full Runtime Results
Figure 6: Runtime decomposition for IVIPER, MAVIPER, and HYDRAVIPER (with full environment interaction budget) on the multi-agent particle world environments, cooperative navigation and physical deception. Error bars show 95% confidence intervals based on 10 random seeds. Runtime measurement was performed following the methodology in Section 5.2.
Figure 7: Runtime decomposition for IVIPER, MAVIPER, and HYDRAVIPER (with full environment interaction budget) on the traffic signal control environments, Cologne corridor and Ingolstadt corridor. Error bars show 95% confidence intervals based on 10 random seeds. Runtime measurement was performed following the methodology in Section 5.2.
Appendix E Full Hyperparameter Sensitivity Results
In Figure 8, Figure 9, Figure 10, and Figure 11, we show the performance of HYDRAVIPER in the low budget setting as two hyperparameters — DT depth and the UCB scaling constant $c$ — are varied for the cooperative navigation (CN), physical deception (PD), Cologne corridor (CC), and Ingolstadt corridor (IC) environments. We use 500 training/1 500 validation rollouts for CN and PD, and 100 training/100 validation rollouts for CC and IC.
Figure 8: Sensitivity of HYDRAVIPER to two hyperparameters, DT depth and the UCB scaling constant $c$, on the cooperative navigation (CN) environment. Lower rewards are better.
@@ Appendix E Full Hyperparameter Sensitivity Results
Figure 10: Sensitivity of HYDRAVIPER to two hyperparameters, DT depth and the UCB scaling constant $c$, on the Cologne corridor (CC) environment. Lower rewards are better.
@@
Appendix F Full Ablation Results
In Figure 12, we show the performance of HYDRAVIPER in the physical deception (PD, left) and Ingolstadt corridor (IC, right) environments, in an ablation study where each of the algorithm’s components — centralised-$Q$ resampling (C-Q), adaptive training budget allocation (TR-A), and UCB-based validation budget allocation (VR-UCB) — is removed in turn. For the high budget level, we use 5 000 training/5 000 validation rollouts for PD, and 1 000 training/1 000 validation rollouts for IC; for the low budget level, we use 500 training/1 500 validation rollouts for PD, and 100 training/100 validation rollouts for IC.
Figure 12: Ablation on physical deception (PD, left, higher rewards are better) and Ingolstadt corridor (IC, right, lower rewards are better).